# User Simulation for Evaluating Information Access Systems

### v1

@@ 5.2 Choice and Decision Making in Recommender Systems
Recommender systems are meant to help users make better choices through a collaborative process between a user and the system. In order to realistically simulate users’ interactions with recommender systems, it is useful to have a broad understanding of how humans make choices.
@@ 5.2.1 Choice Patterns
  • Consequence-based choice: A different way of thinking about choice is to consider the consequences of choosing a particular option. The recommender system can help, for example, by determining (in a personalized way) when there is a choice to be made, identifying options that the user was not aware of that were available, or highlighting (and possibly warning against) particular consequences that the user might not be familiar with.
@@
The above patterns are presented in a distilled form, but in practice they are often used in combination. For example, a simple attribute-based strategy is employed to narrow down the space of options, then another pattern is used to choose among those options.
@@ 5.3 Mathematical Framework
In order to realistically simulate user behaviour, it is important to realize that the actions users take are not independent events, but are part of their decision-making process, even if they are often studied in isolation. The previous sections discussed model of information seeking and decision making in terms of cognitive processes. The simulation of a user, however, also needs a mathematical framework for modeling decision-making from the user’s perspective, i.e., to represent the current state, the set of available actions, and the corresponding rewards. This is most commonly modeled as a Markov decision process (MDP), which can formally be described by a finite state space $\mathcal{S}$, a finite action set $\mathcal{A}$, a set of transition probabilities $P$, and a reward function $R$. In an MDP, we take discrete time steps. At a given point in time, the simulated user is in state $s\in\mathcal{S}$, and by executing action $a\in\mathcal{A}$, they transition into a new state $s^{\prime}$ according to the transition probability $P(s^{\prime}|s,a)$. The Markov property ensures that this transition depends only on the current state and action, without needing to consider the entire history of states and actions, which simplifies modeling and reduces computational complexity. The reward function $R(a,s)$ represents the direct rewards, i.e., costs or contributions, associated with the decision of taking action $a$ at state $s$. Sometimes, the reward is not immediate but can only be received at the end of a sequence of interactions. In an MDP, the goal is to find a policy $\pi$ that maximizes the expected cumulative reward over time.
@@
There are some differences, however, in how MDPs are used in reinforcement learning (RL) and in user simulation. In RL, the main focus revolves around finding an optimal policy [Singh:1999:NIPS]. Designing effective reward functions is crucial to achieve that, while transition probabilities are often observed from an external environment, without modeling them explicitly. Conversely, in user simulation, the reward function can be used to encapsulate the costs and rewards based on observed data (from logs or user studies), while transition probabilities are modeled explicitly based on some model of user behaviour. [^1] Parameters of the user model may hand-crafted and/or learned from data. The policy is viewed as a mapping from states to actions, $\pi(s)\rightarrow a$, which governs the user’s decision making. Even though a rational user behaviour may generally be assumed, the policy does not need to be optimal. Indeed, even humans are unlikely to behave optimally when exposed to a new information access system for the first time, as it takes time to learn how the system works and how it can be utilized most effectively. In user simulation, the policy is also based on an explicit model of user behaviour. Some decisions might be made based on deterministic rules while others vary across users, depending on their personal characteristics. Crucially, for the purposes of simulation-based evaluation, the policy needs to be controllable by the system designer.
@@
We use the MDP framework to systematically organize the simulation of various components of search and recommender systems in Chapter 6. In the context of task-oriented dialogue systems, MDPs have been used for user simulation for a long time; we continue to build on this in Chapter 7.
@@
This chapter focuses on simulating how a user interacts with search engines and recommender systems. These represent the two main modes of information access: pull, where search is initiated by the user, and push, where the system takes initiative and recommends content to the user (cf. Section 1.1). There are several reasons for treating and discussing these two modes together. First, it has been established early on that the two problems are essentially “two sides of the same coin” [Belkin:1992:CACM] and it has also been reiterated more recently that there are theoretical and practical reasons to model the two tasks together [Zamani:2018:DESIRES]. Indeed, the separation between the two tasks is increasingly blurred in modern information service platforms, where search results may be accompanied by related items as recommendations, and recommendations may be filtered based on search queries. Second, there is a lot of common elements when it comes to users interactions with these systems: in both cases, the user needs to examine a ranked list of items, select which ones they find attractive enough to click on, and decide when to stop interacting with the system. There are, of course, differences as well: query formulation is specific to search, while leaving explicit feedback (such as likes or ratings) is a signature characteristic of recommender systems. Overall, there has been considerably more research attention given to user modeling and evaluation for search engines than for recommender systems, which will be reflected in this chapter. For simplicity, we will be using standard terminology from information retrieval throughout most of the chapter. That is, we will be talking about result lists served on search engine result pages (SERPs), consisting of a ranked list of results or snippets that link to documents. However, snippets may be replaced with previews and documents may be replaced with arbitrary information objects (e.g., entities, products, images, videos).
@@
We start in Section 6.1 by presenting models that describe interaction workflows, that is, specify the space of user actions and system responses, and possible transitions between them. From the perspective of the cooperative game framework (cf. Section 4.3), these model define the “moves” that a user and a system can each make. In terms of an MDP framework (cf. Section 5.3), these define the space of actions and the possible transitions between them. Then, we discuss specific user actions: query formulation in Section 6.2, scanning behaviour in Section 6.3, clicks in Section 6.4, effort involved in processing documents in Section 6.5, and stopping in Section 6.6. We present toolkits and resources in Section 6.7 and approaches to validating simulators in Section 6.8. The length and depth of discussion in each of these sections are proportional to the amount of research effort invested in each respective area. We close this chapter in Section 6.9 with a summary, discussing how the various component-level models fit together and identifying where gaps and limitations remain.
@@ 6.1.1 Search Workflows
The Complex Searcher Model, proposed by Maxwell:2015:CIKM and then further updated in [Maxwell:2018:ECIR], combines actions from prior searcher models [Baskaya:2013:CIKM, Thomas:2014:IIiX] and additionally introduces a SERP-level examination decision; see Fig. 6.4. The boxes (A)–(F) on this figure denote sets of interactions: (A) topic examination, which assumes a TREC-like topic description outlining the information need that motivates the search; (B) querying, i.e., the process of generating potential queries to address the underlying information need and deciding which ones to issue; (C) SERP examination, where the searcher obtains an initial impression of the SERP and decides whether it appears promising enough to examine the results in detail or rather it should be abandoned; (D) result summary examination, that is, looking at the result snippets and determining if they look attractive enough to be clicked for further examination; (E) document examination, which involves reading (parts of the) document to assess whether it is relevant to the information need, and “saving” it, if it is; (F) deciding to stop, i.e., making decisions about whether or not to stay on the current SERP and, if not, whether to continue with the search session (and issue further queries) or terminate the session.
@@ 6.1.2 Recommendation Workflow
Figure 6.5: Flowchart of a simple user model for recommender systems.
We are not aware of any workflow models being proposed specifically for recommendation. However, we can imagine that a user would interact with a recommender system according to the workflow depicted in Fig. 6.5. The user examines previews of items in a result list, and they decide for each if they find it attractive enough to click on it. Once they clicked on it (and consumed the item, which is not an action we consider in this simple model), they decide if they want to leave feedback on it. Then, they can either go back to the result list (which may or may not be updated based on feedback left during the session) and continue examining items or stop interacting with the recommender. The respective actions will be discussed in Sections 6.36.6.
@@ 6.2 Simulating Queries
Simulating queries is one of the most difficult aspects of user simulation in IR and as such has attracted considerable attention. We discuss seminal work organized into three main groups: the generation of individual queries, the generation of query sequences based on discriminative power, and query reformulations; see Table 6.1 for an overview. Generally, most methods work by sampling terms from documents/topic language models and often make use of existing test collections, i.e., TREC-like topic definitions (see Fig. 6.6 for an example) and relevance judgments. The generation process may be probabilistic or deterministic in nature. Most methods are static in a sense that they are not affected by what the system does; dynamic query generation, on the other hand, considers previous queries and results returned for those queries. It is worth emphasizing that simulated queries “are synthetic in nature and are not necessarily representative of real user queries” [Jordan:2006:JCDL]. The generated queries can be evaluated by checking whether they reproduce statistical properties of real queries and by comparing the performance of retrieval systems using real vs. simulated queries (cf. Section 6.8).
@@ 6.2.1 Generating Individual Queries
Azzopardi:2007:SIGIR propose generative probabilistic models that generate queries for known item search. It is assumed that the user has a particular, previously seen document in mind, the “known item,” that they want to retrieve. The user thus tries to construct a query by recalling terms of phrases from that document. The generative process is as follows:
  • Initialize an empty query $q=\{\}$.
  • Sample a document $d$ to be the known item with probability $P(d)$.
  • Select the query length $l$ with probability $P(l)$.
  • Repeat $s$ times:
    • Select a term $t_{i}$ from the (unigram) language model of document $d$ with probability $P(t_{i}|\theta_{d})$.
    • Add $t_{i}$ to the query $q$.
  • Record $(q,d)$ as the known-item query-document pair.
Notice that this process generates not only queries but query-document pairs; since there is a single relevant document (the known item), it allows for automatic evaluation with a query set of any size, without needing to collect relevance judgments. A user may also be assumed to recognize the relevant document when the search result contains it and thus click on such a document (see the discussion of click modeling later). To instantiate the above generic process, three probability distributions $P(d)$, $P(l)$, and $P(t_{i}|d)$ need to be defined. These choices characterize the type and style of queries that are generated. The document prior, $P(d)$, may be set to be uniform (i.e., all documents are equally likely to be selected) or non-uniform (favoring more important documents, e.g., based on document inlink count). The query length distribution, $P(l)$, is approximated with a Poisson distribution (with the mean set to the average length of actual queries). The process of sampling query terms from a document is modeled as a mixture between sampling from a particular document ($d$) and from the collection ($C$):
$P(t_{i}|\theta_{d})=(1-\lambda)P(t_{i}|d)+\lambda(t_{i}|C)~{}.$ (6.1)
The value of the mixture parameter $\lambda$ is inversely proportional to the user’s recollection of the original document. (Based on a set of real queries, Azzopardi:2007:SIGIR use $\lambda=0.2$.) The likelihood of selecting a term from the document, $P(t_{i}|d)$, defines the term selection strategy that is used. For example, users may be assumed to (1) select terms randomly, (2) be more likely to pick frequent terms, or (3) try to choose terms that discriminate the document from other documents in the collection. It is worth noting that all these strategies sample query terms independently. In order to better resemble real queries, the dependencies between query terms would also need to be captured by conditioning term selection on previous query terms, i.e., $P(t_{i}|t_{i-1},\dots,t_{1},\theta_{d})$.
Azzopardi:2009:SIGIR extend the above process for the generation of ad hoc queries. Here, it is assumed that a topic $T=(q_{0},R)$ is given as input, which is defined by a seed query $q_{0}$ and the corresponding set of relevant documents $R$. Azzopardi:2009:SIGIR uses TREC collections, with topic titles as seed queries. Queries of a given length are generated by repeatedly sampling terms according to the following (unigram) query language model:
$P(t_{i}|\theta_{q})=(1-\lambda)P(t_{i}|\theta_{T})+\lambda(t_{i}|C)~{},$ (6.2)
where the first term corresponds to the selection of terms that are on topic, while the second term corresponds to sampling from the background collection (i.e., noise). To estimate the topic model $P(t_{i}|\theta_{T})$, possible strategies include selecting frequent or discriminative terms from the set $R$ of relevant documents, or conditioning term selection on the seed query $q_{0}$.

### v2

@@ 5.2 Choice and Decision Making in Recommender Systems
Recommender systems are meant to help users make better choices through a collaborative process between a user and the system. In order to realistically simulate users’ interactions with recommender systems, it is useful to have a broad understanding of how humans make choices. Indeed, a realistic user simulator should attempt to simulate how humans make such choices in various scenarios when interacting with the options provided by a system.
@@ 5.2.1 Choice Patterns
  • Consequence-based choice: A different way of thinking about choice is to consider the consequences of choosing a particular option. The recommender system can help, for example, by determining (in a personalized way) when there is a choice to be made, identifying available options that the user was not aware of, or highlighting (and possibly warning against) particular consequences that the user might not be familiar with.
@@
The above patterns are presented in a distilled form, but in practice they are often used in combination. For example, a simple attribute-based strategy is employed to narrow down the space of options, then another pattern is used to choose among those options. While the six ASPECT choice patterns were originally formulated to explain individual choices, they can also be applied in “scenarios in which there is a need to support the choice of a group of persons” [Jameson:2022:bookchapter].
@@ 5.3 Mathematical Framework
In order to realistically simulate user behaviour, it is important to realize that the actions users take are not independent events. While they are often studied in isolation, these actions are part of a larger decision-making process. Earlier in this chapter we have explored the cognitive side of this process—how users make choices when seeking information or interacting with recommended items. Now, let us introduce a mathematical framework that helps us formally model decision-making from the user’s point of view: the Markov Decision Process (MDP).
Imagine a user interacting with an information access system such as a search engine. At any given moment, they are in a specific state ($s\in\mathcal{S}$, e.g., typing a query, scanning the result list, clicking on a result snippet). They have choices, i.e., a set of actions ($a\in\mathcal{A}$) they can take (e.g., submit a query, click on a result, leave the search engine). Each action leads them to a new state ($s^{\prime}$), like following a path. An MDP helps us model this journey and is made up of:
  • States ($\mathcal{S}$): The different situations a user can be in.
  • Actions ($\mathcal{A}$): The choices available to the user in each state.
  • Transition probabilities ($P(s^{\prime}|s,a)$): The likelihood of moving from one state ($s$) to another ($s^{\prime}$) after taking a specific action ($a$).
  • Rewards ($R(a,s)$): The immediate benefit (or cost) the user gets from taking a particular action ($a$) in a given state ($s$). This could be, e.g., finding the information they seek or experiencing frustration. Sometimes, the reward is not immediate but can only be received at the end of a sequence of interactions. For example, successfully purchasing an item or abandoning the system after a series of failed attempts.
In an MDP, we take discrete time steps. The Markov property, a fundamental assumption of this framework, states that the next state ($s^{\prime}$) the system transitions to depends solely on the current state ($s$) and the action taken ($a$). In other words, the system has no “memory” of its past states and actions. This simplifies modeling and reduces computational complexity.
In the MDP framework, a policy ($\pi:\mathcal{S}\rightarrow\mathcal{A}$) is a function that determines which action to take in each state ($\pi(s)\rightarrow a$). That is, the policy guides the simulated user’s behaviour, dictating which action to choose when faced with different scenarios within the system.
MDPs are commonly used for modeling systems that make sequences of decisions to maximize rewards, such as in robotics or game playing. In these applications, the MDP framework enables reinforcement learning (RL), where an agent (that controls the system’s decision-making) learns to interact with an environment through trial and error. The agent takes actions, observes the resulting states and rewards, and adjusts its policy iteratively to achieve the highest possible cumulative reward over time.
@@
Although both reinforcement learning and user simulation utilize MDPs, there are important differences arising from their distinct objectives. In RL, the main focus revolves around finding an optimal policy [Singh:1999:NIPS]. Designing effective reward functions is crucial to achieve that, while transition probabilities are often observed from an external environment, without necessarily modeling them explicitly, though in model-based RL, predictive models for state transitions and actions may be used [Moerland:2023:Survey]. Conversely, in user simulation, the reward function can be used to encapsulate the costs and rewards based on observed data (from logs or user studies), while transition probabilities are modeled explicitly based on some model of user behaviour. [^1] Parameters of the user model may be hand-crafted and/or learned from data. Even though a rational user behaviour may generally be assumed, the policy, governing the user’s decision making, does not need to be optimal. Indeed, humans are unlikely to behave optimally when exposed to a new information access system for the first time, as it takes time to learn how the system works and how it can be utilized most effectively. In user simulation, the policy is also based on an explicit model of user behaviour. Some decisions might be made based on deterministic rules while others vary across users, depending on their personal characteristics. Crucially, for the purposes of simulation-based evaluation, the policy needs to be controllable by the system designer.
@@
One benefit of framing the user simulation problem as to create a user simulation agent based on MDP is the natural connection with various learning algorithms such as RL and imitation learning [Hussein:2017:CSUR, Labhishetty:2023:thesis]. It also allows systematic discussion of all components in a user simulator. We use the MDP framework to systematically organize the simulation of various components of search and recommender systems in Chapter 6. In the context of task-oriented dialogue systems, MDPs have been used for user simulation for a long time; we continue to build on this in Chapter 7.
@@
This chapter focuses on simulating how a user interacts with search engines and recommender systems. These represent the two main modes of information access: pull, where search is initiated by the user, and push, where the system takes initiative and recommends content to the user (cf. Section 1.1). There are several reasons for treating and discussing these two modes together. First, it has been established early on that the two problems are essentially “two sides of the same coin” [Belkin:1992:CACM] and it has also been reiterated more recently that there are theoretical and practical reasons to model the two tasks together [Zamani:2018:DESIRES]. Indeed, the separation between the two tasks is increasingly blurred in modern information service platforms, where search results may be accompanied by related items as recommendations, and recommendations may be filtered based on search queries. Second, there is a lot of common elements when it comes to how a user interacts with these systems: in both cases, the user needs to examine a ranked list of items, select which ones they find attractive enough to click on, and decide when to stop interacting with the system. There are, of course, differences as well: query formulation is specific to search, while leaving explicit feedback (such as likes or ratings) is a signature characteristic of recommender systems. Overall, there has been considerably more research attention given to user modeling and evaluation for search engines than for recommender systems, which will be reflected in this chapter. For simplicity, we will be using standard terminology from information retrieval throughout most of the chapter. That is, we will be talking about result lists served on search engine result pages (SERPs), consisting of a ranked list of results or snippets that link to documents. However, snippets may be replaced with previews and documents may be replaced with arbitrary information objects (e.g., entities, products, images, videos).
@@
We start in Section 6.1 by presenting models that describe interaction workflows, specifying the space of user actions and system responses, and possible transitions between them. From the perspective of the cooperative game framework [Zhai:2016:IEEE] (cf. Section 4.3), these models define the “moves” that a user and a system can each make. In terms of a Markov Decision Process (MDP) framework (cf. Section 5.3), these define the space of actions and the possible transitions between them. Then, we discuss specific user actions: query formulation in Section 6.2, scanning behaviour in Section 6.3, clicks in Section 6.4, effort involved in processing documents in Section 6.5, and stopping in Section 6.6. The length and depth of discussion in each of these sections are proportional to the amount of research effort invested in each respective area. We close this chapter in Section 6.7 with a summary, discussing how the various component-level models fit together and identifying where gaps and limitations remain.
@@ 6.1.1 Search Workflows
The Complex Searcher Model, proposed by Maxwell:2015:CIKM and then further updated in [Maxwell:2018:ECIR], combines actions from prior searcher models [Baskaya:2013:CIKM, Thomas:2014:IIiX] and additionally introduces a SERP-level examination decision; see Fig. 6.4. The boxes (A)–(F) on this figure denote sets of interactions: (A) topic examination, which assumes a TREC-like topic description outlining the information need that motivates the search; [^2] For ordinary users, without TREC-like topic descriptions, this step corresponds to reflecting on their information need. (B) querying, i.e., the process of generating potential queries to address the underlying information need and deciding which ones to issue; (C) SERP examination, where the searcher obtains an initial impression of the SERP and decides whether it appears promising enough to examine the results in detail or rather it should be abandoned; (D) result summary examination, that is, looking at the result snippets and determining if they look attractive enough to be clicked for further examination; (E) document examination, which involves reading (parts of the) document to assess whether it is relevant to the information need, and “saving” it, if it is; (F) deciding to stop, i.e., making decisions about whether or not to stay on the current SERP and, if not, whether to continue with the search session (and issue further queries) or terminate the session.
@@ 6.1.2 Recommendation Workflow
Figure 6.5: Flowchart of a basic user model for recommender systems that present recommendations as a ranked list.
Figure 6.6: Flowchart of an advanced user model for recommender systems, corresponding to carousel-based interfaces with multiple ranked lists (rows). The main components are: (A) row examination, (B) item examination, (C) feedback, and (D) stopping decisions.
We are not aware of any workflow models being proposed specifically for recommendation, though elements of a workflow pertaining to users’ scanning and clicking behaviour in carousel-based user interfaces are described in [Rahdari:2024:TORS]. In the most basic setting, where recommendations are presented as a ranked list, we can imagine that a user would interact with a recommender system according to the following simplified workflow, depicted in Fig. 6.5. The user examines previews of items in a result list, and they decide for each if they find it attractive enough to click on it. Once they clicked on it (and consumed the item, which is not an action we consider in this simple model), they decide if they want to leave feedback on it. Then, they can either go back to the result list (which may or may not be updated based on feedback left during the session) and continue examining items or stop interacting with the recommender.
Clearly, the above workflow is largely incomplete as it does not consider how modern recommender systems typically present results using carousel-based interfaces (cf. Section 6.3.3). There, the results are not presented as a single ranked list but rather as several ranked lists in multiple rows, each labeled with a category or topic. Inspired by the work of Rahdari:2024:TORS, we present an advanced workflow in Fig. 6.6 corresponding to carousel-based interfaces. Each carousel contains a ranked list of items, which we refer to simply as a row. It is assumed that the user examines rows in sequential order. If they find a row attractive, they will examine the items contained in the row, also in sequential order, and can click on them and leave feedback the same way as in the basic model. They can decide to stop examining items within a row and, then, to continue to the next row or stop.
This workflow might be further extended based on additional functionalities offered by modern recommender systems, such as exploratory browsing by showing additional recommendations based on the current item being viewed. These additional recommendations can appear in various forms, such as related items that are similar or complementary (e.g., in an e-commerce setting, under “similar products” and “others also bought” sections), encouraging a deeper exploration of available items. However, this is beyond the scope of this book, and the development of more complex workflows that correspond to modern recommender system UIs remains an open area for future research.
The various user actions depicted in Figs. 6.5 and 6.6 will be discussed in Sections 6.36.6.
6.1.3 Multiple Workflows
It is worth noting that multiple workflows, including both homogeneous workflows (e.g., multiple search engine workflows or multiple recommendation workflows) and heterogeneous workflows (e.g., search workflow combined with recommendation workflow), can generally be combined in a mixed-mode information access system, potentially generating infinitely many different variations of workflows. From a user simulation perspective, it means that for different systems with different workflows, we generally need to model user actions differently since the user actions could vary significantly. Another factor is the context or state in the MDP framework used for modeling a user, which clearly depends on the specific workflow. A general way of modeling interactions with any interface is to use the Interface Card Model [Zhang:2015:SIGIR], where a user is modeled with three component probabilistic models, i.e., probability of actions to be taken on any given interface, the cost of performing each action, and the utility of performing each action. How this model can be used to support a general simulation-based evaluation framework was discussed earlier in Chapter 4.
@@ 6.2 Simulating Queries
The problem of simulating queries has attracted considerable attention in IR. We discuss seminal work organized into three main groups: the generation of individual queries, the generation of query sequences based on discriminative power, and query reformulations; see Table 6.1 for an overview. Generally, most methods work by sampling terms from documents/topic language models and often make use of existing test collections, i.e., TREC-like topic definitions (see Fig. 6.7 for an example) and relevance judgments. The generation process may be probabilistic or deterministic in nature. Most methods are static in a sense that they are not affected by what the system does; dynamic query generation, on the other hand, considers previous queries and results returned for those queries. It is worth emphasizing that simulated queries “are synthetic in nature and are not necessarily representative of real user queries” [Jordan:2006:JCDL]. For example, Breuer:2022:ECIR show that is is possible to reproduce statistical properties of real queries and preserve the relative system orderings using simulated queries, but the query terms are very different from those used in real queries.
@@ 6.2.1 Generating Individual Queries
Azzopardi:2007:SIGIR propose generative probabilistic models that generate queries for known item search. It is assumed that the user has a particular, previously seen document in mind, the “known item,” that they want to retrieve. The user thus tries to construct a query by recalling terms of phrases from that document. The process involves sampling a document $d$ to be the known item with probability $P(d)$, selecting a query length $l$ with probability $P(l)$, and then creating a query by sampling terms $t_{1},\dots,t_{l}$ from the (unigram) language model of document $d$ with probability $P(t_{i}|\theta_{d})$. The probability distributions $P(d)$, $P(l)$, and $P(t_{i}|d)$ characterize the type and style of queries that are generated.
Notice that this process generates not only queries but query-document pairs; since there is a single relevant document (the known item), it allows for automatic evaluation with a query set of any size, without needing to collect relevance judgments. A user may also be assumed to recognize the relevant document when the search result contains it and thus click on such a document (see the discussion of click modeling later).
Azzopardi:2009:SIGIR extend the above process for the generation of ad hoc queries. Here, it is assumed that a topic $T=(q_{0},R)$ is given as input, which is defined by a seed query $q_{0}$ (i.e., TREC topic titles) and the corresponding set of relevant documents $R$. Queries of a given length are generated by repeatedly sampling terms according to a two-component mixture query language model, where the first term ($P(t_{i}|\theta_{T})$) corresponds to the selection of terms that are on topic, while the second term (P$(t_{i}|C)$) corresponds to sampling from the background collection (i.e., noise). To estimate the topic model, possible strategies include selecting frequent or discriminative terms from the set $R$ of relevant documents, or conditioning term selection on the seed query $q_{0}$.